### v2

# Discovery of Critical Thresholds in Mixed Exposures and Estimation of Policy Intervention Effects using Targeted Learning

Abstract
Traditional regulations of chemical exposure tend to focus on single exposures, overlooking the potential amplified toxicity due to multiple concurrent exposures. We are interested in understanding the average outcome if exposures were limited to fall under a multivariate threshold. Because threshold levels are often unknown a priori, we provide an algorithm that finds exposure threshold levels where the expected outcome is maximized or minimized. Because both identifying thresholds and estimating policy effects on the same data would lead to overfitting bias, we also provide a data-adaptive estimation framework, which allows for both threshold discovery and policy estimation. Simulation studies show asymptotic convergence to the optimal exposure region and to the true effect of an intervention. We demonstrate how our method identifies true interactions in a public synthetic mixture data set. Finally, we applied our method to NHANES data to discover metal exposures that have the most harmful effects on telomere length. We provide an implementation in the CVtreeMLE R package.
Keywords: Targeted Maximum Likelihood Estimation, Mixtures, Interactions, Decision Trees, Ensemble Learning
1 Introduction
Policymakers and health researchers seek to understand how joint exposures impact health outcomes and to find and establish safe exposure thresholds based on concurrent exposure. Traditionally, regulatory frameworks have been oriented towards setting exposure limits based on individual chemicals. However, in the complex environment of real-world settings, individuals are invariably exposed to a plethora of chemicals simultaneously, often referred to as mixed exposures. Such exposures range from air pollutants and endocrine disruptors to pesticides and heavy metals. Landrigan and Goldman [2006], Vrijheid et al. [2009], Kortenkamp [2014], Braun and Gennings [2018], Council [2018].
The complexity of mixed exposures arises when the joint impact of such chemical mixtures deviates from the additive effects of their individual components. Combined exposures could potentially increase health risks beyond the sum of the effects of individual agents NRC [2008]. In contrast, some exposures might counteract the impacts of others. Furthermore, the effects of these combined exposures can vary due to environmental stressors, genetic predispositions, and psychosocial determinants Carlin et al. [2013]. Given this complexity, there is an urgent call for cutting-edge, assumption-lean statistical methodologies suitable for delimitating exposure thresholds for concurrent exposure Safe [1993], Kortenkamp [2007] to help inform safe exposure limits.
In modern causal inference concerning mixed exposures, conventional approaches have relied heavily on parametric models. More often than not, consideration is given to the effect of a singular exposure, typically represented as a coefficient within a generalized linear regression model (GLM). Such piecemeal evaluations fail to capture the collective effect of multiple concurrent exposures, a common scenario in realistic exposure settings Davidson et al. [2009], Kim et al. [2014], Eze et al. [2011]. Although efforts have been made to estimate the joint effects of mixed exposures - examples including the popular Weighted Quantile Sum Regression (WQS) Keil et al. [2019], Bayesian Mixture Modeling De Vocht et al. [2012] and Bayesian Kernel Machine Regression Bobb et al. [2014] - they come with their caveats. Some of these methods harbor strong inherent assumptions, such as the directional homogeneity and linear/additive assumptions in WQS. Bayesian approaches demand prior knowledge or risk human-induced bias due to subjective choice of priors. Flexible methods such as BKMR are often visualized through a variety of dose-response graphs rather than distilled into an intuitive summary statistic Bobb et al. [2014]. To date, none of the methods have been built to efficiently estimate a predefined target parameter, starting from a full-causal model and explicitly stating the assumptions necessary to identify the parameter in the observed data van der Laan and Rose [2011, 2018]. We attempt to bridge this gap.
Ideally, the analyst would have the full multidimensional dose-response relationship, but this is complex due to unrealistic assumptions and nonpathwise differentiability, complicating robust confidence interval construction. To address this, we suggest categorizing joint exposures to assess the impact of specific exposure thresholds on average outcomes. If our aim is to quantify the effect of imposing certain exposure thresholds, a sensible approach is to categorize the joint exposure and compare the average outcome if everyone were forced to have exposures within a region compared to the observed average outcome under observed exposures. Because this simulated policy is based on a region discovered in the exposure space, we call our target parameter the average regional exposure effect (ARE). This approach is helpful because interpretable categories can be defined, such as $(A_{1}<a_{1})\&(A_{2}<a_{2})$, where $a_{i}$ are specific values in $A$ that are of clear interest to policy makers. Identification assumptions are also more transparent in this setting. In this context, we seek to categorize the exposure space and find the regions with the expected minimum (or maximum) population outcome. However, this intervention that effectively forces people to a region is ambiguous. The post-intervention probability density of exposure $A$ is concentrated in the given region, but there are infinitely many different policies that achieve this. Our choice preserves relative natural selection preferences within the region, per covariate stratum, which is a natural option and lends itself to easy estimation with standard tools. However, the thresholding of the exposure space that minimizes/maximizes the population average outcome is not known a priori. Data must be used to determine which regions are relevant. To estimate this oracle region of the exposure space, we develop a recursive partitioning algorithm paired with g-computation to greedily search the exposure space. Once the thresholds have been identified, we obtain the most precise inference possible for the effects using a doubly robust and efficient estimator, targeted maximum likelihood estimation (TMLE).
Using the same data to define thresholds and estimate effects in the outcome minimum region can lead to biased results. We address this problem by splitting the data, estimating thresholds on one part, and estimating counterfactual effects (given the fixed thresholds) on the other. To efficiently use all the data, we repeat this splitting process in a round-robin fashion (K-fold cross-validation). We build this framework on previous work related to data-adaptive parameters Hubbard et al. [2016] and cross-validated targeted minimum loss-based estimation (CV-TMLE) Zheng and van der Laan [2010], our method, called CVtreeMLE, is a novel approach for estimating the joint impact of a mixed exposure by using CV-TMLE that guarantees consistency, efficiency, and multiple robustness despite using highly flexible learners (ensemble machine learning) to estimate a data-adaptive parameter. A crucial challenge in this process is the inconsistency of the data-adaptive parameter. This is especially evident in the case of variable importance methods, which use quantiles of variables and estimates variable importance based on the average treatment effect (ATE), measuring quantiles with the highest expected mean against the lowest. Hubbard et al. [2018] There is potential variability of the parameters identified (quantiles) across different data subsets that are identified as the maximum or minimum in the respective fold. Such variability can complicate the interpretation of these data-adaptive target parameters. Our current work seeks to address this inherent challenge of a mixture of possibly different regions found across the folds for our target parameter. We not only provide pooled estimates common to data-adaptive target parameters for the maximizing/minimizing W-controlled exposure region, which may be an amalgamation of different regions due to variability in findings across the folds, but also introduce a new approach to pool estimates across different exposure sets. For example, if thresholds are found for $(A_{1}<a_{1})\&(A_{2}<a_{2})$ in 6 of 10 folds, because thresholds are found for this exposure set $A_{1}$ and $A_{2}$, we pool these estimates for this exposure set.
We provide implementations of this methodology in our free and open-source software package CVtreeMLE, written in the R language and environment for statistical computing (R Core Team, 2022). The CVtreeMLE software package has undergone a rigorous peer review process by the Journal of Open Source Software, which validates the robust implementation of our innovative methodology for the analysis of mixed exposure McCoy et al. [2023].
This manuscript is structured as follows: Section 2 discusses semiparametric methodology and redefines the population intervention effect as the average regional effect (ARE). Section 3 focuses on ARE estimation and inference, while Section 4 introduces the oracle target parameter. Section 5 delves into cross-estimation, including pooling methods. Simulation studies demonstrating the estimator’s properties are in Section 6. Section 7 applies CVtreeMLE to NIEHS data and compares it with other methods, section 7.2 we apply our method to NHANES data. The final sections (8 and 9) detail the CVtreeMLE software and conclude with a discussion.
2 The Estimation Problem
2.1 Setup and Notation
We consider an observational study with baseline covariates ($W\in\mathbb{R}^{p}$), multiple exposures ($A\in\mathbb{R}^{m}$), and a single timepoint outcome ($Y$). Let $O=(W,A,Y)$ denote the observable data. We assume that there exists a potential outcome function $Y(a)$ (that is, $Y(a)$ is a random variable for each value of $a$) that generates the outcome that would have been obtained for each observation had exposure been forced to the value $A=a$. These potential outcomes are not observed, but the observed outcome $Y$ corresponds to the potential outcome for the observed value $A$ of the exposure, that is, $Y=Y(A)$. We use $P_{0}$ to denote the true data-generating distribution. We assume that our $O_{1},O_{2},...,O_{n}$ are independent and identically distributed (i.i.d.) draws of $O=(W,A,Y)\sim P_{0}$. We decompose the joint density as $p_{Y,A,W}(y,a,w)=p_{Y|A,W}(y,a,w)p_{A|W}(a,w)p_{W}(w)$ and make no assumptions about the forms of these densities. Let $\mu(a,w)=E[Y|A=a,W=w]$ represent the observable (stratum-specific) dose-response curve that can actually be estimated from observable data.
Let $\Psi$ map from the set of possible data-generating distributions to the real numbers. Let $\Psi(P_{0})$, the value of this map at the true, unknown distribution $P_{0}$, denote a target parameter (estimand) of interest. We can view our observed data $(O_{1}\dots O_{n})$ as a (random) probability distribution $P_{n}$ that assigns a probability mass $1/n$ to each observation $O_{i}$. Let $\hat{\Psi}$ mapping from empirical distributions to numbers denote our estimator (algorithm) such that $\hat{\Psi}(P_{n})$ is our estimate from data.
2.2 Defining the Differential Effect Given Regional Exposure
[…]
$=E[E[Y|\mathcal{A}=1,W=w]]$
What we have shown is that this parameter is a population average of some function $Q(\mathcal{A},W)=E[Y|\mathcal{A},W]$ when $\mathcal{A}$ is forced to 1. For any value of $w$, $Q(1,w)$ is a particular convexly weighted average of the causal dose-response curve across the different exposure levels, thus collapsing it down to a single number for each $w$.We want to compare the average outcome if all individuals had exposures within the region $\mathcal{A}$ with the average observed outcome under the observed exposure distribution. This defines our parameter of interest, which we call the average regional effect (ARE): $\psi=E[Q(\mathcal{A}=1,W)]-E[Y]$, representing the difference in the average outcomes if we forced exposure to self-selection within $\mathcal{A}$ versus the observed outcomes. In most applications, it is not known a priori which region $\mathcal{A}$ should be set. For example, we may not know how various chemicals or drugs interact and how to set safe limits for all of them. Therefore, $\mathcal{A}$ itself is in practice something that should be estimated to maximize some objective. That said, for the purposes of establishing our theory, we should first imagine $\mathcal{A}$ as known and fixed. We will then show how we can choose what policy to enact (i.e., choose $\mathcal{A}$) while also unbiasedly estimating its effect. We first discuss the assumption necessary for the statistical quantity estimable from sample data to have causal interpretations, then we move on to discuss how $\mathcal{A}$ can be determined.
2.3 Identification and Causal Assumptions
Our target parameter is defined in terms of the observable distribution and does not represent a causal quantity without additional assumptions. To add a causal interpretation, we define the causal ARE as $\psi^{*}=E[Y(A)|\mathcal{A}=1]-E[Y]$. Standard conditioning arguments show $\psi^{*}=\psi$ when the conditional randomization assumption, $A\perp Y(a)\mid W$ for all $a\in\mathcal{A}$ and positivity: $P(\mathcal{A}=1|W)>0$ for all $w$, are met. The randomization assumption, asserts that the potential outcomes are independent of exposure within covariate-defined strata, indicating that there are no unmeasured confounders. The positivity assumption, $P(\mathcal{A}=1|W)>0$ for all $w$, ensures that all exposure levels and covariate strata are represented in the data, critical for accurate effect estimation. By satisfying these conditions, we can estimate the causal ARE reliably, focusing on obtaining the observable ARE efficiently without additional assumptions such as linearity. Although our identification assumptions may not always hold in all applications, we can at least eliminate model misspecification bias and minimize random variation. Once we have established how to estimate the ARE for a fixed region, we will turn our attention to the problem of finding a good region $\mathcal{A}$ and finally how to do that without incurring selection bias in estimating the ARE for that region.
3 Estimating ARE with TMLE
In the previous sections, we established that the causal ARE is equivalent to the observable ARE $E[E[Y|\mathcal{A}=1,W]-E[Y]]$ under standard identification assumptions. Therefore, to estimate it, all we need to do is 1) create a new binary random variable $\mathcal{A}_{i}=1_{\mathcal{A}}(A_{i})$ and 2) proceed as if we were estimating the observable average counterfactual effect $E[E[Y|\mathcal{A}=1,W]]$ from the observational data structure $(Y,\mathcal{A},W)$ where $\mathcal{A}$ is a simple binary “exposure”.
We can adapt techniques for estimating the ATE $E[E[Y|\mathcal{A}=1,W]]-E[E[Y|\mathcal{A}=0,W]]$ to construct ARE estimators that offer provable unbiasedness (assuming that there is no bias from potential violations of identification assumptions) and efficiency. These estimators are optimized for minimal sampling variance and possess the "doubly robust" property, as highlighted in Zheng and van der Laan [2010], Zivich and Breskin [2021]. Both Augmented Inverse Propensity-Weighting (AIPW) and Targeted Maximum Likelihood Estimation (TMLE) stand out as assumption-free and data-efficient methods when used with machine learning and cross-fitting. Although they often produce similar results, TMLE has been found to perform more efficiently with smaller samples Luque-Fernandez et al. [2018], Smith et al. [2022], van der Laan and Rose [2011, 2018], making it a generally preferable approach. For this reason, we construct a TMLE estimator for the ARE. The TMLE estimator is inspired by the fact that if we knew the true conditional mean $Q(\mathcal{A},W)=E[Y|\mathcal{A},W]$, we could estimate the ARE with the empirical average $\frac{1}{n}\sum_{i}Q(1,W_{i})-Y_{i}$. Of course, we do not know $Q$, but we can estimate it by regressing the outcome $Y$ on the exposure $\mathcal{A}$ and the covariates $W$. However, a detailed mathematical analysis shows that we incur bias if we use our estimate $\hat{Q}$ instead of the truth. This bias might decrease as the sample size increases, but it dominates relative to random variability, making it impossible to establish p-values or confidence intervals. TMLE solves this problem by computing a correction to the regression model $\hat{Q}$ that removes this bias. In other words, “targets” the estimate $\hat{Q}$ to the parameter of interest (here, ARE). We use targeted learning to de-bias our initial predictions of the outcome if all individuals were exposed to exposure levels within the region. The process is as follows: 1. Use cross-validated ensembles of machine learning algorithms (a “super learner”) to generate estimates of the conditional means of treatment: $\hat{g}(\mathcal{A}=a,W)\approx P(\mathcal{A}=a|W)$ (i.e. propensity score) and outcome: $\hat{Q}(\mathcal{A},W)\approx E[Y|\mathcal{A},W]$. 2. Regress $Y$ (scaled to $[0,1]$) onto the “clever covariate” $H_{i}=\frac{\pi(\mathcal{A}_{i})}{\hat{g}(1,W_{i})}$ using a logistic regression with weights $\text{logit}(\hat{Q}(\mathcal{A},W))$. Here, $\pi(\mathcal{A}_{i})$ is the marginal probability of being in the region. The (rescaled) output of this is our target regression model $\hat{Q}^{*}$. That is, using $\epsilon$ from this least favorable model, we debias our predictions for $\hat{Q}(\mathcal{A}=1,W)$. 3. Compute the plug-in estimate using the targeted model: $\hat{\psi}=\frac{1}{n}\sum_{i}\hat{Q}^{*}(1,W_{i})-Y_{i}$. An estimated standard error for $\hat{\psi}$ is given by: $\hat{\sigma}^{2}=\frac{1}{n^{2}}\sum_{i}\left[\left(\frac{\pi(\mathcal{A}_{i})}{\hat{g}(1,W_{i})}\right)\bigg{(}Y_{i}-\hat{Q}^{*}(1,W_{i})\bigg{)}+\bigg{(}\hat{Q}^{*}(1,W_{i})-Y_{i}-\hat{\psi}\bigg{)}-\bigg{(}Y_{i}-E[Y]\bigg{)}\right]^{2}$.
With corresponding 95% confidence interval $\hat{\psi}\pm 1.96\hat{\sigma}$. Here, two terms before the minus are the efficient influence function (EIF) for the parameter $E[E[Y|\mathcal{A}=1,W]]$ and the third term is the EIF for the mean of the random variable $Y$. Explaining why the targeting step takes the form of a logistic regression and how the estimated standard error is derived is beyond the scope of this work. van der Laan and Rubin [2006], van der Laan and Rose [2011, 2018] offer explanations targeted to audiences with varying levels of mathematical sophistication. To obtain these estimates, we need only to specify the ensemble of machine learning algorithms used to estimate the propensity and initial outcome regressions $\hat{g}$ and $\hat{Q}$. Theoretical guarantees hold as long as a sufficiently rich library is chosen. To estimate the ARE, we must also specify the region $\mathcal{A}$ so that we can compute our binary “exposure” variable. The issue of course is that we have been treating $\mathcal{A}$ as a known region, whereas in many applications the important question is determining what guidelines to impose in the first place. This is the focus of the next section.
4 Finding a Good Exposure Region
[…]
4.1 G-Computation Decision Tree
We introduce a novel decision tree algorithm developed to approximate the oracle parameter. This parameter, denoted as $\underset{\mathcal{A}}{\mathrm{argmin}}\ E[E[Y|\mathcal{A}=1,W]]$, represents the region $\mathcal{A}$ that minimizes the expected outcome for a given set of covariates $W$. Our approach is tailored to meet several key criteria: computational efficiency, interpretability, and implicit regularization to avoid overly granular and statistically insignificant regions. The core of our method is a regression tree that optimizes a simple plug-in estimator for the policy effect ($E[Y|\tilde{A}=1,W]$) of a given region, based on a Random Forest or a similar straightforward model. The decision tree is constructed to search for the minimum/maximum region using this plug-in estimator, greedily evaluating splits in the exposure variables $A$ and selecting the region that maximizes/minimizes the estimated average outcome in the newly formed regions.
In the partitioning criterion phase, the algorithm explores all binary splits across the range of each exposure variable to identify the optimal partitioning strategy. Consider a scenario where we have five exposure variables, A1 through A5. The algorithm examines the unique values for each exposure variable. For example, for A1, it evaluates the potential splits at A1 < 1, A1 < 2, and so forth, up to A1 < m, where m represents the maximum observed value of A1. At each of these split points, the algorithm constructs a binary indicator defining the region of interest and employs the plug-in estimator. This estimator is trained to assess the mean outcome under the hypothetical situation where all individuals are ‘forced’ into that region. The algorithm’s objective at each split point is to minimize/maximize the average predicted outcome within the newly formed child nodes, relative to the parent node. This process of finding the optimal split is applied to each of the exposure variables. For example, if the algorithm identifies that splitting at A5 < 5 yields the minimum average outcome, this split becomes the root node of our decision tree. Subsequently, the algorithm partitions the data into two segments based on the chosen split, resulting in a left node (A5 $<$ 5) and a right node (A5 $\geq$ 5). The recursive process is then applied to each of these data regions demarcated by the node, iterating through all exposure variables again to find further splits that minimize the average outcome. For example, within the left node, the algorithm might test combinations such as A5 $<$ 5 &\&& A1 $<$ 2, etc., each time training the plug-in estimator and evaluating the mean outcome for the defined region. This recursive partitioning continues until two stopping criteria are met: it either reaches the pre-specified maximum depth of the tree, or it arrives at a point where no further split can yield an average outcome lower/greater than that of the parent node.
5 K-fold Cross-Estimation
As discussed, the identification of interaction regions and the estimation of ARE requires a data partitioning strategy that mitigates bias. We employ a K-fold cross-estimation technique that ensures the asymptotic properties of our estimators without reliance on additional assumptions. This involves dividing the data set into complementary estimation ($P_{n_{k}}$) and parameter-generating ($P_{n_{-k}}$) samples, the latter being utilized to derive exposure regions and train nuisance parameters essential for the target maximum likelihood estimate (TMLE) update. For each fold, we determine the exposure thresholds using $P_{n_{-k}}$ and, with the same sample, train our nuisance estimators $g_{n}$ and $Q_{n}$. Subsequently, these estimators are applied to $P_{n_{k}}$ to obtain unbiased ARE estimates within the fold.
5.1 Pooled TMLE
Upon completion of the cross-estimation procedure, a pooled TMLE update provides a summary measure of the oracle region target parameter across k-folds. Specifically, we stack the predictions for each nuisance parameter from the estimation samples and run a pooled TMLE update on the cumulative initial estimates. The resulting average is then used for parameter estimation.
Our substitution estimator, denoted by $\bar{Q}_{n}$, approximates the true conditional mean $\bar{Q}_{0}$ by plugging in the empirical distribution for each observation. The estimator is operationalized via a Super Learner algorithm followed by TMLE, and its cross-validated variant is expressed as $\Psi(Q_{P_{n_{k}}}^{\star})=\frac{1}{V}\sum_{v=1}^{V}\{\bar{Q}_{n_{-k}}^{\star}(\mathcal{A}_{n_{-k}}=1,W_{v})-Y_{v}\}$. Where $Q_{n_{-k}}^{\star}$ is the TMLE updated expectation of the outcome if everyone were exposed to levels of exposure within the data-adaptively determined region. The cross-estimation approach not only utilizes the full dataset for variance estimation, resulting in tighter confidence intervals, but also enables the derivation of fold-specific and pooled estimates.
That is, this pooled parameter represents the ARE for the determined maxmizing/minimizing region in the exposure space. Of course, the actual regions determined in folds can vary, and thus this acts as an omnibus test, which can be used to evaluate fold-specific results. A significant pooled TMLE ARE for the oracle region can be interpreted as ’There is a region in the exposure space where the expected outcome is significantly different compared to the average outcome.’ The analyst can then evaluate the fold-specific results to determine what regions make up this oracle estimate. Of course, interpretability of the pooled oracle estimate is reliant on regions in the same sets of exposures that are found across the folds. Otherwise, this pooled estimate allows the analyst to understand if there is generally "signal" in the data for a region or set of regions that collectively differ in expected outcomes compared to the average. This dual presentation allows for the evaluation of variability across folds, offering insight into the stability of the pooled ARE. In particular, when exposure regions $\mathcal{A}$ exhibit high variability between folds, fold-specific results provide an essential interpretive counterbalance to the pooled estimates.
We also provide TMLE updated estimates for the region for the same sets of exposures. For example, if a region was found using thresholds for exposures A1 and A3 in 7 out of 10 folds, we pool the results for these folds for this exposure set. Of course, because 3 folds did not include these exposures for the oracle region, we need to incorporate an inferred null estimate and weigh our pooled estimate incorporating these null results. We discuss this weighted pooling in the next section.
5.2 Pooled TMLE Incorporating Null Results
It is possible that the region with the minimum or maximum expected outcome is not found in every fold. This makes the pooled result difficult to interpret, as these estimates do not incorporate information from folds with null results. As such, in refining our approach to pooled TMLE estimation, our objective is to integrate both the non-null estimates from the estimation samples and the null values. We propose a weighted average approach that combines our original pooled TMLE estimate with the implied null value, which we define as 0. Specifically, let $n_{0}$ denote the number of folds where the null is implied, and let $n_{1}$ represent the number of folds with non-null estimates. The variance of the "null" estimate, $var(\Psi_{\text{null}})$, is estimated using the variance of the original pooled estimates, $var(\hat{\Psi}_{\text{pooled}})$, as $var(\Psi_{\text{null}})=\left(var(\hat{\Psi}_{\text{pooled}})\right)\cdot\frac{n_{1}}{n_{0}}$. We then calculate the weighted average of the original pooled estimate and the null value: $\hat{\Psi}_{\text{pooled}}^{*}=\frac{w_{1}\cdot\hat{\Psi}_{\text{pooled}}+w_{0}\cdot\Psi_{\text{null}}}{w_{0}+w_{1}}$, where $w_{1}=\frac{1}{var(\hat{\Psi}_{\text{pooled}})}$ and $w_{0}=\frac{1}{var(\Psi_{\text{null}})}$. The variance of this new pooled estimate is given by $var(\hat{\Psi}_{\text{pooled}}^{*})=\frac{1}{w_{1}+w_{0}}$. This approach ensures that our pooled TMLE estimate reflects both the explicit estimates from folds where the maximizing/minimizing region was selected and the implicit null values from the remaining folds. In the case where the oracle region in all folds is contained in the same exposure set, we can average the rules to create an interpretable rule for the pooled TMLE estimate. For example, given three partitions like $X_{1}<2$ $X_{2}>5$, $X_{1}<2.3$ $X_{2}>5.2$, and $X_{1}<1.9$ $X_{2}>5.3$, the average rule becomes $X_{1}<2.06(1.9,2.3)$ $X_{2}>5.16(5,5.3)$. The numbers in parentheses are the lower and upper bounds of the threshold for the oracle region, which give an estimate of uncertainty for the thresholds. Overall, interpretability of the estimate can be straightforward if there is consistency in where the oracle region is in the exposure space across folds. If the same exposures contain the oracle region across all the folds, the analyst can simply interpret the pooled TMLE ARE as the estimate for the average oracle region. However, if there is variability, the pooled TMLE acts as a sanity check for any set of regions, leading to a significant difference in the average outcome. If so, the analyst can then look at the k-fold specific results, which also have valid inference, and the pooled estimates within each exposure set while also reporting the number of folds a specific exposure set contained the oracle region.
6 Simulations
In this section, we demonstrate using simulations that our approach identifies the correct exposure region that maximizes/minimizes conditional mean and estimates the correct PIE target parameter built into a DGP for this region.
6.1 Data-Generating Processes
Because a two-dimensional exposure space is easier to visualize and describe compared to higher-dimensional spaces, we created a squared dose-response relationship between two exposure variables where an interaction occurs between the exposures when each meets a particular threshold value. Specific outcome values were generated for each subspace of the mixture based on split points $\mathcal{A}_{d}$, but there exists a region with the maximum outcome (the truth that we want). The goal is to determine whether our adaptive target parameter is targeting the region that maximizes/minimizes the conditional mean outcome for the given sample and to evaluate how CVtreeMLE approaches this desired oracle parameter (the true PIE for this region) as the sample size increases. To achieve this goal, we construct a data-generating process (DGP) explained below.
6.1.1 Two-Dimensional Exposure Simulations
This DGP has the following characteristics, $O=(W,A,Y)$. $W$ are three baseline covariates, $W_{1}\sim\mathcal{N}(\mu=37,\sigma=3),W_{2}\sim\mathcal{N}(\mu=20,\sigma=1),W_{3}\sim\mathcal{B}(\mu=0.5)$. Here $\mathcal{B}$ is a Bernoulli distribution, and $\mathcal{N}$ is normal. These distributions and values were chosen to represent a study with covariates for age, BMI, and sex. Our generated exposures were also created to represent two chemical exposures quantized into five discrete levels. The values and range of the outcomes were chosen to represent common environmental health outcomes, such as telomere length or epigenetic expression. We are interested in sampling observations into a 2-dimensional exposure grid. We use discrete exposures to facilitate calculation of the ground truth effects without numerical integration. Here, a $5\times 5$ grid is based on combinations of two discrete exposure levels with values 1-5. We want the number of observations in each of these cells to be affected by covariates. To do this, we define a conditional categorical distribution $P\{(A_{1},A_{2})=(a_{1},a_{2})|W=w\}$ and sample from it. $P\{(A_{1},A_{2})=(a_{k},a_{l})|W\}=\frac{e^{W^{\top}\beta_{k,l}}}{1+\sum_{k,l}e^{W^{\top}\beta_{k,l}}}$. Here, the $\beta$’s attached to each covariate were drawn from a normal distribution with means 0.3, 0.4, 0.5 and 0.5, respectively, all with a standard deviation of 2. This then gives us 25 unique exposure regions with densities dependent on the covariates. We then want to assign an outcome to each of these regions based on the main effects and interactions between exposures. We generate the outcome as: $Y=0.2A_{1}^{2}+0.5A_{1}A_{2}+0.5A_{2}^{2}+0.2*W_{1}+0.4*W_{3}+\mathcal{N}(0,0.1)$. This indicates that there is a slightly weaker squared effect for $A_{2}$ relative to $A_{1}$ and a strong interaction between the exposures and confounding due to $W_{1}$ and $W_{2}$.
Computing Ground Truth
The fact that our exposures are discrete in this simulation allows us to easily compute the ground truth mean outcome for any region $\mathcal{A}$ because we can explicitly compute the conditional mean function $Q_{\mathcal{A}}$
[…]
$=\frac{\sum_{a\in\mathcal{A}}\mu(a,w)p_{A|W}(a,w)}{\sum_{a\in\mathcal{A}}p_{A|W}(a)}$
Therefore, to approximate the ARE of a region $\mathcal{A}$ to arbitrary precision we compute $\text{ARE}(\mathcal{A})=\frac{1}{b}\sum_{i}^{b}Q_{\mathcal{A}}(1,W_{i})-Y_{i}$ using a large $b$. We repeat this for each of the $(6^{2})^{2}$ possible rectangular regions to determine which has the optimal ARE.
6.2 Evaluating Performance
The following steps describe how our simulation was tested to determine 1. asymptotic convergence to the true oracle region used in the DGP, 2. convergence to the true ARE based on this oracle region and 3. convergence to the true data-adaptive ARE, that is, CVtreeMLE’s ability to correctly estimate the ARE if the data-determined rule was applied to the population. We do this by first generating a random sample from the DGP of size $n$ which is broken into $K$ equal size estimation samples of size $n_{k}=n/K$ with the corresponding parameter that generates samples of size $n-n/K$. At each iteration, the parameter generating fold is used to define the region via our algorithm and is used to estimate the necessary nuisance estimators $\hat{Q}$ and $\hat{g}$. The estimation fold is used to get our TMLE updated estimate, and then we do this for all folds. For an iteration, we output the ARE estimates given the pooled TMLE and the k-fold specific TMLE. We compare the estimated and ground truth AREs for each region identified in each fold to estimate the data-adaptive bias. We also compare estimated AREs and regions with oracle equivalents.
For each iteration, we calculate metrics for bias, variance, MSE, CI coverage, and confusion table metrics for the true minimal region compared to the estimated region. We also calculate the bias for the estimated mean in the region found in our recursive partitioning algorithm to find the minimum region. For each type of estimate (pooled TMLE, k-fold specific TMLE estimates, and inverse variance, if applicable), we have bias when comparing our estimate to 1. The ARE based on the true region in the DGP that minimizes the outcome mean 2. The ARE when the data-adaptively determined region is applied to the population. Therefore, when comparing to the true "oracle" region ARE we have $\psi^{0}_{\text{pooled tmle MSE}}$: This is the MSE of the pooled TMLE ARE compared to the ground-truth ARE for the true oracle region built into the DGP which minimizes the mean adjusted outcome and $\psi^{0}_{\text{mean v-fold tmle MSE}}$: This is the MSE of the mean k-fold specific AREs compared to the ground-truth ARE for the true region built into the DGP.
We are also interested in the ARE if the data-adaptive-determined region, the region estimated to minimize the adjusted outcome in the sample data, were applied to $P_{0}$ the true population. Therefore, there are also MSE estimates for $\psi^{DA}_{\text{pooled tmle MSE}}$: This is the MSE of the pooled TMLE ARE compared to the ARE of the average region across the folds applied to $P_{0}$ and $\psi^{DA}_{\text{mean v-fold tmle MSE}}$: This is the MSE of the mean k-fold specific AREs compared to the mean ARE when all the k-fold specific rules are applied to $P_{0}$.
For each ARE estimate we calculate the confidence interval coverage of the true ARE parameter given the oracle region and the ARE given the data-adaptively determined region applied to $P_{0}$. For the TMLE pooled estimates, these are lower and upper confidence intervals based on the pooled influence curve. For the k-fold-specific coverage, we take the mean lower and upper bounds. In each case, we check to see if the ground-truth rule ARE and the data-adaptive rule ARE are within the interval.
Lastly, we compare the data-adaptively identified region (the region determined by our recursive partitioning algorithm to be the minimal region amongst all regions) with the ground-truth region (built into the DGP) using the confusion table metrics for true positive, true negative, false positive, and false negative to determine whether, as the sample size increases, we converge to the true region. Similarly, we compare the expected outcome for the minimum region calculated in our recursive partitioning algorithm to the true minimum outcome to ensure that our estimates are converging to the truth. These performance metrics were calculated in each iteration, where 100 iterations were performed for each sample size $n=$ (200, 350, 500, 750, 1000, 1500, 2000, 3000, 5000). It was ensured that for each data sample, at least one observation existed in the ground truth region to ensure that confusion table estimates could be calculated. CVtreeMLE was run with a 10-fold CV with default learner stacks for each nuisance parameter and data-adaptive parameter.
6.3 Default Estimators
As discussed, CVtreeMLE needs estimators for $Q=E(Y|A,W)$ and $g=P(A|W)$. CVtreeMLE has built in default algorithms to be used in a Super Learner van der Laan Mark et al. [2007] that are fast and flexible. These include random forest, general linear models, elastic net, and xgboost. These are used to create Super Learners for $Q_{n}$ and $g_{n}$. Users can pass in their own libraries for these nuisance and data-adaptive parameters. For these simulations, we use these default estimators in each Super Learner. Our partitioning algorithm to discover the region $\mathcal{A}$ was set to have a maximum depth of 3 and a minimum number of observations per leaf to 10. In this simulation we are interested in finding the region with maximum outcome. The package has a "min_max" parameter which we set to "max" to find the region with maximum expected mean outcome.
6.4 Results
6.4.1 CVtreeMLE Algorithm Identifies the Oracle Region with Minimum Expected Outcome
First, we describe the results to identify the true region built into the DGP. It is obviously necessary that this converges to the truth as the sample size increases in order for the estimates $\psi^{0}$ to be asymptotically unbiased. Overall, we find that our tree algorithm identifies the true region in the DGP and therefore provides results that have a minimum outcome for treatment policies. Our recursive algorithm shows very good identification of the true outcome minimizing region even in low sample sizes. At sample size 300 the average false positives were 0.053, at 500, 0.0005, and at 750 onward 0. The average true positives were 100% across all sample sizes. Our recursive partitioning algorithm correctly identifies the interaction between both exposures as the minimizing region (as opposed to marginal impacts) in 100% of folds in all sample sizes. As such, we discuss the pooled TMLE results which are the same as the inverse variance results, since in no fold were there null results. Figure 1 shows the bias in comparing the estimated average in the estimated minimum region of the decision tree to that of the true oracle average in the oracle minimizing region. From this figure it can be seen that we have asymptotic convergence to the true minimal average at a rate faster than $\sqrt{n}$ for this simulation.
Figure 1: Average Bias from G-computation Plug-in from the Decision Tree Algorithm
6.4.2 CVtreeMLE Unbiasedly Estimates the Data-Adaptive Parameter
Looking at the MSE for the ARE estimate between the adaptive region discovered in the data and the ARE when this region is applied to $P_{0}$, TMLE unbiasedly estimates the adaptive parameter of the data at root rates $n$ with an average of 95% CI coverage. Below, Figure 2 A shows the data-adaptive rule ARE MSE ($\psi^{DA}_{\text{pooled tmle MSE}}$, $\psi^{DA}_{\text{mean k-fold tmle MSE}}$). Given that the pooled TMLE is the average estimates across folds after the pooled update, we would expect these estimates to be fairly close, which they are. Coverage by sample size was: 300 = 98%, 750 = 94%, 2000 = 95%, 5000 = 95 %.
Figure 2: Data-Adaptive and Oracle Parameter MSE
6.4.3 CVtreeMLE Unbiasedly Estimates the Oracle Target Parameter
Now, we compare the estimates with the true ARE given the oracle region in the DGP. Figure 2 B show the MSE for this comparison in our simulation. As can be seen, both decrease faster than the root $n$ rate for both the fold-specific and pooled estimates. Based on these simulations, CVtreeMLE unbiasedly estimates the oracle target parameter at root $n$ rates. Coverage by sample size was: 300 = 97%, 750 = 95%, 2000 = 94%, 5000 = 95 %.
7 Applications
7.1 NIEHS Synthetic Mixtures
The NIEHS synthetic mixture data (found here in github) are a commonly used data set to evaluate the performance of statistical methods for mixtures. These synthetic data can be considered as the results of a prospective cohort study. The outcome cannot cause exposures (as might occur in a cross-sectional study). Correlations between exposure variables can be thought of as caused by common sources or modes of exposure. The nuisance variable $W$ can be assumed to be a potential confounder and not a collider. There are 7 exposures ($A_{1}-A_{7}$) that have a complicated dependency structure with a biologically based dose response function based on endocrine disruption. For details, the synthetic data key on the github page for data set 1 (used here) gives a description of how the data were generated. Generally, there are two groups of exposures ($A_{1},A_{2},A_{3}$ and $A_{5},A_{6}$). Therefore, the correlations within these clusters are high. $A_{1},A_{2},A_{7}$ contribute positively to the outcome; $A_{4},A_{5}$ contribute negatively; $A_{3}$ and $A_{6}$ do not have an impact on the outcome, making rejecting these variables difficult given their correlations with members of the group. This correlation and effects structure is biologically plausible because different congeners of a group of compounds (e.g., PCBs) may be highly correlated but have different biological effects. There are various agonistic and antagonistic interactions in exposures. A breakdown of the variables sets and their relationships are: $A_{1},A_{2}$ both increase Y by concetration addition, $A_{1}-A_{4},A_{2}-A_{4},A_{1}-A_{4},A_{2}-A_{4}$ competetive antagonism, $A_{1}-A_{7},A_{2}-A_{7}$ supra-additive or synergistic, $A_{4},A_{5}$ both addivively decrease Y, $A_{4}-A_{7},A_{5}-A_{7}$ unusual antagonism. Given these toxicological interactions, we can expect certain statistical interactions determined as cut-points for sets of variables from CVtreeMLE. Since we are interested in one region that minimizes the mean outcome under some restrictions, we expect to find a region that includes lower levels for an exposure with an agonistic relationship or perhaps an interaction between an agonistic and antagonistic exposure. The NIEHS data set has 500 observations and 9 variables. $A$ is a binary confounder. Of course, in this data there is no ground truth, like in the above simulations, but we can gauge CVtreeMLE’s performance by determining if the expected variable sets are included in the rule which defines the minimizing region. We apply CVtreeMLE to this NIEHS synthetic data using a 20-fold CV and the default stacks of estimators used in the Super Learner for all parameters. We parallelize over the cross-validation to test computational run-time on a newer personal machine an analyst might be using. We show results when searching for both the minimizing region and the maximizing region. The minimum number of observations in a region was set to 20.
In 20 of 20 folds, a region in $X_{2}$ was found to have the lowest adjusted outcome. The average rule was found to be $X_{2}<=0.382$. This threshold varied between 0.38 and 0.39 a very low partition variability. The ARE for this region was found to be -4.39 (-6.37 - -2.42), which means that if all individuals were exposed to levels of $X_{2}$ within this region, the result would be 4.47 units less than the observed average. This effect was significant (p < 0.001). The fact that this region was found in 100% of folds indicates a robust result. This result can be interpreted as, in order to have the largest impact to reduce the outcome, this exposure should be adjusted levels below this threshold. For the maximum region with at least 20 observations, we found that X1 has the highest mean outcome at thresholds greater than 3.3 (3.06 - 3.49), restricting observations below this threshold had a borderline significant effect of 8.21 (-3.40 - 19.84). This indicates that if all individuals were forced to this region the outcome would be 8.21 larger than observed average. A threshold for this exposure X1 was found in 100% of folds.
The dose-response curves for these exposures in these synthetic data are provided on the github page. In each case, these results are reasonable given the marginal dose response relationships. Finally, given the antagonistic relationships built in, we might expect that the region with the lowest mean may be an interaction between an agonist and antagonist which we have not detected because of restricting observations in the region, when we reduce the min observations to 10 for the regions, we detect this minimizing region (Table 1). These results give an example of variability, since this minimizing region was found in 90% of the folds that included exposures X2 and X5 and one fold for X2 and X7. The TMLE result of the pooled oracle region was -7.17 (-29.37 - 15.04, p = 0.53) which indicates that the ARE averaged between these two regions is -7.17. Table 1 results provide an example of the results of inverse variance as the detected region changed slightly between folds. From this example, the analyst would report a consistent antagonistic relationship found between exposures X2 and X5 in 90% of the folds with a significant ARE.
ARE
SE
Lower CI
Upper CI
P-value
Average Rule
Prop.
-6.802
1.218
-9.190
-4.413
0.001
X2 <=0.384(0.38,0.39) & X5 >=1.122(0.87,1.23)
0.90
-0.286
6.393
-12.816
12.244
0.96
X2 <= 0.38 & X7 <= 0.44
0.05
Table 1: NIEHS Results: Inverse Variance Estimates for the Minimizing Interaction
7.1.1 Comparison to Existing Methods
Quantile g-computation Keil et al. [2019], prevalent in environmental epidemiology for mixture analysis, estimates the effects of uniformly increasing exposures by one quantile, based on linear model assumptions. This method quantizes mixture components, summing the linear model’s coefficients to form a summary measure ($\Psi$) for joint impact assessment. However, it inherently assumes additive, monotonic exposure-response relationships, overlooking complex, potentially nonlinear interactions typical in mixtures, like in endocrine disrupting compounds. Consequently, this method might not accurately capture the nuanced dynamics of mixed exposures, especially when interactions vary with other variable levels.
We run quantile g-computation on the NIEHS data using 4 quantiles with no interactions to investigate results using this model. The size of the scaled effect (positive direction, sum of positive coefficients) was 6.28 and included $A_{1},A_{2},A_{3},A_{7}$ and the scaled effect size (negative direction, sum of negative coefficients) was -3.68 and included $A_{4},A_{5},A_{6}$. Compared to NIEHS ground truth, $A_{3},A_{6}$ are incorrectly included in these estimates. However, the positive and negative associations for the other variables are correct. Next, because we expect interactions to exist in the mixture data, we would like to assess for them but the question is which interaction terms to include? Our best guess is to include interaction terms for all exposures. We do this and show results in Table 2.
Estimate Std. Error Lower CI Upper CI Pr($>$||||t||||)
(Intercept) 21.29 1.58 18.19 24.39 0.00
psi1 0.02 1.62 -3.16 3.20 0.99
psi2 0.59 0.67 -0.71 1.90 0.37
Table 2: Quantile G-Computation Interaction Results from NIEHS Synthetic Data
In Table 2 $\Psi_{1}$ is the summary measure for the main effects and $\Psi_{2}$ for interactions. As can be seen, when including all interactions, neither of the estimates are significant. Of course, this is to be expected given the number of parameters in the model and the sample size $n=500$. However, moving forward with interaction assessment is difficult; if we were to assess for all 2-way interaction of 7 exposures, the number of sets is 21 and with 3-way interactions is 35. We would have to run these many models and then correct for multiple testing. Hopefully, this example shows why mixtures are inherently a data-adaptive problem and why popular methods such as this, although succinct and interpretable, fall short even in a simple synthetic data set.
7.2 NHANES Data
Chemical and metal exposures are linked to changes in telomere length, a key factor in cellular aging and disease. Our applied study explores the collective influence of metal exposure on LTL, aiming to: 1) Present CVtreeMLE application results on real-world data, 2) Share data and processing code via the open-source CVtreeMLE package, and 3) Identify the most harmful exposure levels on telomere length. We utilize data from the NHANES 1999-2002, encompassing demographic information, disease history, nine urinary metals, and LTL for 2510 participants. Confounders include age, gender, race, and lifestyle factors. Metals measured are barium, cadmium, cobalt, cesium, molybdenum, lead, antimony, thallium, and tungsten. Our analysis with CVtreeMLE involves a 10-fold CV and seeks regions where metal exposure most reduces LTL. Data and more information are available in the CVtreeMLE package for further research.
Table 3: Inverse Variance Pooled TMLE Results for NHANES POP Exposures
ARE
SD E.
Lower CI
Upper CI
P-value
Average Rule
Prop.
-0.051
0.26
-0.566
0.465
0.85
cadmium $>$ 2.147 (2.04,2.32)
0.30
-0.173
0.965
-2.065
1.719
0.86
cadmium $>$ 1.93 & cesium $>$ 7.4
0.10
-0.154
0.390
-0.919
0.611
0.7
cadmium $\geq$ 1.888 (1.81,2.03) & tungsten $\leq$0.098 (0.09,0.11)
0.60
The ARE for the oracle parameter across the folds was found to be -0.22 (-0.98 - 0.01), p = 0.06. This indicates that for a region in the metal mixture space that minimizes telomere length, if all individuals were exposed to this region the telomere length would reduce by 0.22. Given this omnibus result, we now review Table 3 which shows the inverse pooled TMLE ARE results for the minimal region found in the folds that make up this oracle region. Here, we see that the detected region that minimizes the average telomere length varies slightly across the folds; however, all regions include cadmium. The region that is the most consistent (60 % of folds) is when cadmium is at or above levels of 1.89 (1.81 - 2.03) and tungsten is at or below levels of 0.098 (0.09 - 0.11). If our NHANES population were forced to be exposed to these levels of cadmium and tungsten, we expect 0.154 base pairs to be shorter than the length of the telomere under observed exposures to cadmium and tungsten. This result was found in 6 of 10 of our folds (running a higher CV can make this more stable). The rules presented are for the average partition across the folds; the numbers in brackets show the lower and upper bounds of the partitions for the respective variable in the rule across the folds and thus can be used to assess stability. We simply imputed missing values at the mean; a more thorough review of better preprocessing methods may improve consistency of results, this is simply an applied example.
In general, in this NHANES example, we show that in real-world data, CVtreeMLE can answer questions about 1. In a high-dimensional mixed metal exposure space, what is the region that minimizes telomere length? and 2. If we were to force the population to have exposures at those levels, what is the expected outcome compared to our observed population outcomes? In this case, we are not answering what the safe levels are, but what the worse levels are, since a shorter telomere length is indicative of faster biological aging.
Our study highlights the need for further research on the health impacts of cadmium and tungsten interactions, particularly on telomere length. Cadmium, a known toxic metal, is linked to health issues like oxidative stress and shorter telomeres. Tungsten, less toxic and less studied, might interact with cadmium in the body, possibly mitigating its harmful effects. The precise nature of this interaction and its effect on cellular aging biomarkers, like telomeres, remains unclear and warrants deeper investigation. This evidence suggests that, when mixed metals are regulated with the aim of reducing the effects on telomere length, the regulation of cadmium and tungsten should be prioritized. Additionally, if populations are exposed to lower levels of tungsten, the regulatory level of cadmium should be lower.
8 Software
The CVtreeMLE R package, is a pivotal tool for using our proposed method. It offers comprehensive documentation, including a vignette on semiparametric theory, real-life examples, and comparative analyses. Emphasizing reproducibility, the package includes applications with NIEHS synthetic data and NHANES mixed metal exposure data. Notably, it supports efficient computational performance, suitable for standard personal computers and scalable to high-performance computing for simulations. The package, available on GitHub, exemplifies the push towards accessible, well-documented, and robust statistical software in scientific research.
9 Discussion
In our study, we introduce a novel approach for analyzing mixed exposures, centering on the ’oracle target parameter’ to determine specific exposure regions’ outcomes. This method employs a decision tree and plug-in g-computation for estimations of these regions. Our goal is to efficiently estimate intervention effects on the maximizing or minimizing region in each fold, utilizing targeted learning within a cross-validation framework to estimate the ARE for this region. This research has significant public policy implications, particularly in identifying safe exposure levels and assessing the impacts of various exposure combinations. A limitation is the inverse variance weighting used to pool within exposure set specific regions to aid in interpretability and gain greater power for these estimates rather than relying on k-fold specific results. Future work will focus on more flexible decision tree development and new methods for incorporating null fold findings with testing. We demonstrate the method’s robustness through simulations and real-world data, culminating in the development of the CVtreeMLE R package for wider application.